In a cross between individuals with the following genotypes for 5 hypothetical genes: AaBbCCDdee x aaBBccDdEe; what is the probability that that one of the progeny would have a dominant phenotype (assume classical dominant/recessive pattern of inheritance for all 5 genes.?

1/16

3/16

1/32

3/32
Answer #1

AaBbCCDdee x aaBBccDdEe

Now we will calculate the probability of getting dominant phenotype for each gene,

for A gene,

A a
a Aa aa
a Aa aa

So, the probability = 1/2.

for B gene,

B b
B BB Bb
B BB Bb

So, the probability = 1.

for C gene,

C C
c Cc Cc
c Cc Cc

So, the probability = 1.

for gene D,

D d
D DD Dd
d Dd dd

So, the probability = 3/4.

for gene E,

e e
E Ee Ee
e ee ee

So, the probability = 1/2.

So, the probability that that one of the progeny would have a dominant phenotype = 1/2 * 1 * 1 * 3/4 * 1/2 = 3/16.